# Multiple Imputation of Hierarchical Nonlinear Time Series Data with an Application to School Enrollment Data

### v1

@@ 3.2.2 Sequential Decomposition of Joint Model
The MINTS method uses a variation of the joint modeling approach to multiple imputation. In the usual joint modeling approach, the complete data $(\mathbf{X},\mathbf{Y})$ is assumed to follow a multivariate joint distribution and imputed values are sampled from the joint posterior predictive distribution of the missing data given the observed data. One downside to this approach is the difficulty of specifying a joint distribution for all variables used in the imputation model. A common general-purpose choice for the joint distribution is the multivariate normal distribution, which has been found to still perform reasonably for imputation even when the assumption of normality is violated (Schafer, 1997a; Schafer and Olsen, 1998). However, if there are nonlinear relationships between variables in the imputation model, joint modeling assuming multivariate normality is unable to incorporate those relationships.
A more flexible approach to multiple imputation is the fully conditional specification (FCS) approach, also known as imputation via chained equations (van Buuren et al., 2006) and sequential regression multivariate imputation (Raghunathan et al., 2001), among other names. FCS does not explicitly specify a joint distribution for the variables with missing data. Instead, univariate conditional distributions are specified for each variable with missing data given all the other variables in the imputation model. Missing values are imputed by iteratively sampling from the univariate conditional distributions until convergence for all imputation model parameters is reached. The conditional distributions can take any form and can accommodate nonlinear relationships between variables. However, this level of flexibility can lead to cases where the univariate conditional distributions are not compatible. Although simulation studies show that FCS can result in reasonable imputations even when the conditional distributions are not compatible (e.g. van Buuren et al., 2006), ultimately FCS does not guarantee that the iterative conditional algorithm will converge to a proper joint distribution.
We use an alternative to the usual joint modeling approach that combines desirable features of joint modeling and FCS through a sequential decomposition of the joint model. The sequential decomposition approach for joint modeling was first proposed by Lipsitz and Ibrahim (1996) and has been extended by Ibrahim et al. (1999, 2002); Lee and Mitra (2016); Xu et al. (2016); Lüdtke et al. (2020); among others. The joint distribution of the variables with missing data is decomposed into a sequence of univariate conditional distributions. One possible decomposition of the joint distribution for $(\mathbf{X},\mathbf{Y})$ is
@@ 3.2.3 Model Specification
where $TN_{[X_{low},X_{up}]}$ refers to the truncated normal distribution with lower bound $X_{low}$ and upper bound $X_{up}$. These boundaries can vary with $(c,t)$, but the dependency is suppressed in the notation.
@@
where $TN_{[Y_{low},Y_{up}]}$ refers to the truncated normal distribution with lower bound $Y_{low}$ and upper bound $Y_{up}$. These boundaries can vary with $(c,t)$, but the dependency is suppressed in the notation.
@@ 3.2.4 Prior Distributions
The truncation is such that $X_{c,0}\in[X_{0,low},X_{0,up}]$ and $Y_{c,0}\in[Y_{0,low},Y_{0,up}]$ where the boundaries can vary with $c$ but this dependency is suppressed in the notation. The control parameters $\bm{\mu}_{early}$ and $\bm{\Sigma}_{early}$ are shared across all $c$.
Priors were chosen to be conjugate and diffuse for most parameters. An informative prior was used for the autoregressive parameter $\rho$ in the imputation model for $\mathbf{Y}|\mathbf{X}$ to reflect the prior belief that $\mathbf{Y}$ is generally increasing over time for the motivating enrollment data, but should generally be specified based on the data being imputed. The values for all hyperparameters should ideally be determined using prior expert knowledge. However, in missing data problems where a general-purpose imputation method is used, the imputer generally does not possess such knowledge for the choice of hyperparameters. In the absence of expert information, we instead propose an algorithm for specifying diffuse priors dictated by data-based control parameters, details of which are in the Appendix. The prior distributions were chosen to be sufficiently diffuse such that the data-based control parameters do not overwhelm the posterior inference following the philosophy of Edwards et al. (1963). We note that use of the data-based algorithm to specify prior distributions results in an approximate posterior distribution rather than a fully Bayesian posterior distribution.
@@ 3.2.5 A-splines
The nonlinear functions $f$ and $h$ in Equation 3 are estimated through spline regression using the complete cases in $(\mathbf{X},\mathbf{Y})$. To estimate $f$, the model $Y_{c,t}=f(X_{c,t})+\varepsilon^{f}_{c,t}$ is fit with Gaussian errors using a B-spline of degree 1. The residuals from the estimation of $f$ are then used to estimate $h$ by fitting the model $|Y_{c,t}-f(X_{c,t})|=h(X_{c,t})+\varepsilon^{h}_{c,t}$ with Gaussian errors and using a B-spline of degree 1. After estimation, $f$ is truncated to have range $[Y_{low},Y_{up}]$ and $h$ is truncated to have range $[\epsilon,\infty)$, where $\epsilon$ is a small positive value. The number and placement of knots for the B-splines used for $f$ and $h$ are selected using a method called adaptive splines, or A-splines (Goepp et al., 2018). A-splines automates the selection of knots using an iterative penalized likelihood approach and is implemented in the R package “aspline” (Goepp, 2022).
@@ 3.3.1 Model Estimation
Estimation of the imputation model parameters occurs simultaneously with estimation of the missing values in a similar fashion to the data augmentation algorithm of Tanner and Wong (1987), with the MCMC algorithm resulting in samples from the joint posterior distribution of the imputation model parameters and the missing values given the observed data. At each iteration of the MCMC algorithm, two steps are iterated until convergence is reached. First, values of the imputation model parameters are drawn from their posterior distributions given the observed data and the most recent estimates of the missing values. Second, estimates for the missing values are drawn from their posterior distributions given the observed data and the previously drawn imputation model parameters. Let $\bm{\theta}_{X}=(\bm{\gamma},\sigma_{X}^{2},\mathbf{X}_{0},\mu_{drift},\sigma^{2}_{drift})$ and $\bm{\theta}_{Y}=(\bm{\alpha},\beta,\rho,\sigma_{Y}^{2},\mathbf{Y}_{0},\mu_{0},\sigma^{2}_{0})$ denote the parameters of the models for $\mathbf{X}$ and $\mathbf{Y}|\mathbf{X}$, respectively. The general approach of the MCMC algorithm proceeds as follows for iterations $i=1,...,n_{iter}$:
@@
The total number of iterations $n_{iter}$ used in the estimation phase is determined based on convergence diagnostics such as inspection of trace plots and evaluation of the diagnostics of Raftery and Lewis (1996) and Gelman and Rubin (1992). Complete details of the MCMC algorithm can be found in the Appendix.
@@ 4 Simulation Study
We conducted a simulation study to evaluate how well the MINTS method performs for estimation of analysis models that are uncongenial to the imputation model. We refer to this validation exercise as “analysis model validation.”
Analysis model validation was conducted for a simulated data set where a nonlinear relationship was simulated between variables. We considered nine experiments corresponding to three rates of simulated missingness and three missing data mechanisms. Each experiment was replicated $N_{rep}=1000$ times and the average performance across replications for MINTS was compared with the performance of existing multiple imputation methods for hierarchical time series data. Details of the simulation study using the nonlinear simulated data are presented in this section, while details of an analogous simulation study using linear simulated data are available in the Appendix.
@@ 4.1 Data Generation
where $TN_{[0,\text{min}(X_{c,t},60)]}$ refers to the truncated normal distribution with support $[0,\text{min}(X_{c,t},60)]$. $\mathbf{X}$ and $\mathbf{Y}$ were constructed to have a generally monotonically increasing relationship similar to the relationship observed for the enrollment data.
@@ 4.2 Analysis Model
Additional validation results for a random intercept analysis model can be found in the Appendix.
@@ 4.3 Analysis Model Validation Procedure
Data was simulated as missing following the three missing data mechanisms of Rubin (1976): Missing Completely At Random (MCAR), Missing At Random (MAR), and Missing Not at Random (MNAR). For variable $\mathbf{X}$, MCAR occurs when $P(\mathbf{R}^{X}|\mathbf{X})=P(\mathbf{R}^{X})$. Data was simulated as missing under MCAR by assuming that each observation has the same probability of being missing. MAR occurs when $P(\mathbf{R}^{X}|\mathbf{X})=P(\mathbf{R}^{X}|\mathbf{X}_{obs})$. Data was simulated as missing under MAR by assuming that observations in earlier years are more likely to be missing. Finally, MNAR occurs when the probability of being missing depends on both the missing and the observed data and $P(\mathbf{R}^{X}|\mathbf{X})$ cannot be simplified further. Data was simulated as missing under MNAR by assuming that the probability of being missing depends on the observed values. All methods compared in the simulation study assume that the missing data mechanism is at least MAR, so the simulations using MNAR act as a sensitivity analysis. Details of the MAR and MNAR implementations can be found in the Appendix.
For each missing data mechanism, data was simulated as missing at the 10%, 40%, and 80% rates. Each combination of missing data mechanism and rate was implemented simultaneously for $\mathbf{X}$ and $\mathbf{Y}$ and defines an experiment. For example, the MCAR 10% experiment corresponds to the setting where 10% of $\mathbf{X}$ was simulated as missing under MCAR and 10% of $\mathbf{Y}$ was simulated as missing under MCAR. We note that while 80% is a high rate of missingness, it is similar to the observed rate of missingness for NER in the enrollment data set.
@@
The performance of the point estimates $\bar{Q}_{M}$ was evaluated using the mean absolute error (MAE), calculated as $\frac{1}{N_{rep}}\sum_{r}|\bar{Q}_{M}-Q|$ where the sum is taken over all replications within each experiment. The performance of the 95% confidence intervals for the pooled estimates $\bar{Q}_{M}$ was evaluated by using the mean coverage across all replications within each experiment, calculated as the proportion of intervals that contained the true value. We also evaluated the mean fraction of Fisher information about $Q$ that is missing due to nonresponse, which we abbreviate as FMI for Fraction of Missing Information. In each replication of each experiment, FMI is estimated as
@@
The mean FMI across all replications within each experiment enables assessment of the amount of information about $Q$ that is lost due to the presence of missing data (Schafer, 1997a; Savalei and Rhemtulla, 2012). We note that with $M=40$ imputations, the estimates of FMI may be noisy, so in this case FMI should only be interpreted as an exploratory diagnostic (Bodner, 2008; Enders, 2010).

### v2

@@ 3.2.1 Assumptions
Finally, we assume that the auxiliary variable is observed at least once for each country for which we are interested in imputing the variable of interest. That is, both $\mathbf{X}$ and $\mathbf{Y}$ can contain missing values for any country-year, but we assume that each country has at least one year in which $X_{c,t}$ is observed. This reflects the motivating setting for the proposed imputation method, where the auxiliary variable is more likely to be observed in each country than the variable of interest due to being easier to measure.
@@ 3.2.2 Sequential Decomposition of Joint Model
The MINTS method uses a variation of the joint modeling approach to multiple imputation. In the usual joint modeling approach, the complete data $(\mathbf{X},\mathbf{Y})$ is assumed to follow a multivariate joint distribution and imputed values are sampled from the joint posterior predictive distribution of the missing data given the observed data. One downside to this approach is the difficulty of specifying a joint distribution for all variables used in the imputation model. A common general-purpose choice for the joint distribution is the multivariate normal distribution, which has been found to still perform reasonably well for imputation even when the assumption of normality is violated (Schafer (1997a); Schafer and Olsen (1998)). However, if there are nonlinear relationships between variables in the imputation model, joint modeling assuming multivariate normality is unable to incorporate those relationships.
A more flexible approach to multiple imputation is the fully conditional specification (FCS) approach, also known as imputation via chained equations (van Buuren et al. (2006)) and sequential regression multivariate imputation (Raghunathan et al. (2001)), among other names. FCS does not explicitly specify a joint distribution for the variables with missing data. Instead, univariate conditional distributions are specified for each variable with missing data given all other variables in the imputation model. Missing values are imputed by iteratively sampling from the univariate conditional distributions until convergence for all imputation model parameters is reached. The conditional distributions can take any form and can accommodate nonlinear relationships between variables. However, this level of flexibility can lead to cases where the univariate conditional distributions are not compatible. Although simulation studies show that FCS can result in reasonable imputations even when the conditional distributions are not compatible (e.g. van Buuren et al. (2006)), ultimately FCS does not guarantee that the iterative conditional algorithm will converge to a proper joint distribution (Li, Yu and Rubin (2012)).
We use an alternative to the usual joint modeling approach that combines desirable features of joint modeling and FCS through a sequential decomposition of the joint model. The sequential decomposition approach for joint modeling was first proposed by Lipsitz and Ibrahim (1996) and has been extended by Ibrahim, Lipsitz and Chen (1999); Ibrahim, Chen and Lipsitz (2002); Lee and Mitra (2016); Xu, Daniels and Winterstein (2016); Lüdtke, Robitzsch and West (2020); among others. The joint distribution of the variables with missing data is decomposed into a sequence of univariate conditional distributions. One possible decomposition of the joint distribution for $(\mathbf{X},\mathbf{Y})$ is
@@ 3.2.3 Model Specification
$\gamma_{c}\sim$ $N(\mu_{drift},\sigma_{drift}^{2}),$
where $TN_{[X_{low},X_{up}]}$ refers to the truncated normal distribution with lower bound $X_{low}$ and upper bound $X_{up}$. These boundaries should be specified based on substantive knowledge about the range of $\mathbf{X}$ and can vary with $(c,t)$, but the dependency is suppressed in the notation. The country-specific drift terms $\gamma_{c}$ follow a world-level distribution where $\mu_{drift}$ and $\sigma_{drift}^{2}$ are world-level hyperparameters.
@@
$\alpha_{c}\sim$ $N(\mu_{0},\sigma_{0}^{2}),$
where $TN_{[Y_{low},Y_{up}]}$ refers to the truncated normal distribution with lower bound $Y_{low}$ and upper bound $Y_{up}$. These boundaries should be specified based on substantive knowledge about the range of $\mathbf{Y}$ and can vary with $(c,t)$, but the dependency is suppressed in the notation. The country-specific intercepts $\alpha_{c}$ follow a world-level distribution where $\mu_{0}$ and $\sigma_{0}^{2}$ are world-level hyperparameters.
@@ 3.2.4 Prior Distributions
The truncation is such that $X_{c,0}\in[X_{0,low},X_{0,up}]$ and $Y_{c,0}\in[Y_{0,low},Y_{0,up}]$. These boundaries should be specified based on substantive knowledge about the ranges of $\mathbf{X}$ and $\mathbf{Y}$ and can vary with $c$, but this dependency is suppressed in the notation. The control parameters $\bm{\mu}_{early}$ and $\bm{\Sigma}_{early}$ are shared across all $c$.
Priors were chosen to be conjugate and diffuse for most parameters. An informative prior was used for the autoregressive parameter $\rho$ in the imputation model for $\mathbf{Y}|\mathbf{X}$ to reflect the prior belief that $\mathbf{Y}$ is generally increasing over time for the motivating school enrollment data, but should generally be specified based on the data being imputed. As the spline term $\beta f(X_{c,t})$ and the AR(1) term $\rho Y_{c,t-1}$ are both on the scale of $Y_{c,t}$, the parameters $\beta$ and $\rho$ can be interpreted as weights that represent the relative importance of the spline term and the AR(1) term. The prior distribution for $\beta$ was chosen to be on the same order of magnitude as $\rho$ to reflect this interpretation.
The values for all hyperparameters should ideally be determined using prior expert knowledge. In the absence of sufficient expertise to specify the prior distributions directly, researchers creating multiple imputations are still likely to have non-expert knowledge of the substantive application that can be used to determine the appropriate order of magnitude for the parameters. Using the school enrollment data as an example, researchers may not have the expertise to specify the prior distributions for $\mu_{drift}$ and $\sigma^{2}_{drift}$ directly. However, they may still have vague prior knowledge that the country-specific drift terms for GER should generally be positive and that increases in school enrollment happen slowly over time for most countries.
As an approximate quantification of this type of vague prior knowledge, we propose an algorithm for specifying diffuse priors dictated by data-based control parameters in the absence of prior expert knowledge. The data-based algorithm enables semi-automated specification of prior distributions using sample statistics calculated from the observed data. The control parameters for the prior distributions of the $\mathbf{X}$ model ($\delta_{X},\nu_{drift},\zeta_{drift},\text{and }\delta_{drift}$) are estimated using summary statistics for the observed first differences $X_{c,t}-X_{c,t-1}$ in $\mathbf{X}_{obs}$. Similarly, the control parameters for the prior distributions of the $\mathbf{Y}|\mathbf{X}$ model ($\delta_{Y},\zeta_{0},\text{and }\delta_{0}$) are estimated using summary statistics for the observed first differences $Y_{c,t}-Y_{c,t-1}$ in $\mathbf{Y}_{obs}$. The control parameters $\bm{\mu}_{early}$ and $\bm{\Sigma}_{early}$ for the prior distributions of $(Y_{c,0},X_{c,0})$ are estimated using summary statistics for $\mathbf{Y}$ and $\mathbf{X}$ from a subset of the data that corresponds to “early” values of $t$, where the range of $t$ that is considered “early” is determined using substantive knowledge of the application. Further details of the algorithm can be found in Section 1 of the Supplementary Material (Liu and Raftery (2025)), and an example illustrating how the algorithm is used to specify prior distributions for the school enrollment data can be found in Section 5.4. We note that use of a data-based algorithm to specify prior distributions results in an approximate posterior distribution rather than a fully Bayesian posterior distribution (Darnieder, 2011).
Although the proposed data-based algorithm is designed to specify prior distributions that are sufficiently diffuse such that the control parameters do not overwhelm the posterior inference following the philosophy of Edwards, Lindman and Savage (1963), the prior distributions resulting from the algorithm should not be used blindly. The prior distributions should be checked to ensure they are sensible for the substantive application and do not unnecessarily restrict the parameter space. For the specification of $\bm{\mu}_{early}$ and $\bm{\Sigma}_{early}$, the data-based algorithm assumes that the time series for $\mathbf{X}$ and $\mathbf{Y}$ change smoothly over time and that the unobserved values $(Y_{c,0},X_{c,0})$ are similar to the observed values in the early subset of data. If there is reason to believe these assumptions are violated, the data-based algorithm should be used with caution.
@@ 3.2.5 A-splines
The nonlinear functions $f$ and $h$ in Equation 3 are estimated through spline regression using the complete cases in $(\mathbf{X},\mathbf{Y})$. To estimate $f$, the model $Y_{c,t}=f(X_{c,t})+\varepsilon^{f}_{c,t}$ is fit with Gaussian errors using a B-spline of degree 1. The residuals from the estimation of $f$ are then used to estimate $h$ by fitting the model $|Y_{c,t}-f(X_{c,t})|=h(X_{c,t})+\varepsilon^{h}_{c,t}$ with Gaussian errors and using a B-spline of degree 1. After estimation, $f$ is truncated to have range $[Y_{low},Y_{up}]$ and $h$ is truncated to have range $[\epsilon,\infty)$, where $\epsilon$ is a small positive value. The number and placement of knots for the B-splines used for $f$ and $h$ are selected using a method called adaptive splines, or A-splines (Goepp, Bouaziz and Nuel (2018)). A-splines automates the selection of knots using an iterative penalized likelihood approach and is implemented in the R package “aspline” (Goepp (2022)).
@@ 3.3.1 Model Estimation
Estimation of the imputation model parameters occurs simultaneously with estimation of the missing values in a similar fashion to the data augmentation algorithm of Tanner and Wong (1987), with the MCMC algorithm resulting in samples from the joint posterior distribution of the imputation model parameters and the missing values given the observed data. At each iteration of the MCMC algorithm, two steps are iterated until convergence is reached. First, values of the imputation model parameters are drawn from their posterior distributions given the observed data and the most recent estimates of the missing values. Second, estimates for the missing values are drawn from their posterior distributions given the observed data and the previously drawn imputation model parameters.
The nonlinear functions $f$ and $h$ are estimated before the start of the MCMC algorithm using the complete cases in the data, pooled across countries and times. The same estimates of $f$ and $h$ are used throughout the estimation and imputation phases. $f$ and $h$ are treated as deterministic within the MCMC algorithm, and uncertainty in the estimation of the nonlinear functions is not accounted for.
Let $\bm{\theta}_{X}=(\bm{\gamma},\sigma_{X}^{2},\mathbf{X}_{0},\mu_{drift},\sigma^{2}_{drift})$ and $\bm{\theta}_{Y}=(\bm{\alpha},\beta,\rho,\sigma_{Y}^{2},\mathbf{Y}_{0},\mu_{0},\sigma^{2}_{0})$ denote the parameters of the models for $\mathbf{X}$ and $\mathbf{Y}|\mathbf{X}$, respectively. The general approach of the MCMC algorithm proceeds as follows for iterations $i=1,...,n_{iter}$:
@@
The total number of iterations $n_{iter}$ used in the estimation phase is determined based on convergence diagnostics such as inspection of trace plots and evaluation of the diagnostics of Raftery and Lewis (1996) and Gelman and Rubin (1992). Complete details of the MCMC algorithm can be found in Section 2 of the Supplementary Material (Liu and Raftery (2025)).
@@ 3.3.2 Imputation Procedure
We note that although missing data is imputed for both $\mathbf{X}_{mis}$ and $\mathbf{Y}_{mis}$, the goal of MINTS is to create multiple imputations of $\mathbf{Y}$. The auxiliary variable $\mathbf{X}$ is imputed out of necessity since $\mathbf{X}$ can also have missing data, but the imputed values of $\mathbf{X}$ are not intended to be used in subsequent analyses.
@@ 4 Simulation Study
We conducted a simulation study to evaluate how well the MINTS method performs for estimation of analysis models that are uncongenial to the imputation model, where uncongeniality arises because imputation of missing data and analysis of imputed data are conducted independently and information is not shared between imputation and analysis tasks. We refer to this validation exercise as “analysis model validation.”
Analysis model validation was conducted for a simulated data set where a nonlinear relationship was simulated between variables. We considered nine experiments corresponding to three rates of simulated missingness and three missing data mechanisms. Each experiment was replicated $N_{rep}=1000$ times. The average performance across replications for MINTS was compared with the performance of existing multiple imputation methods for hierarchical time series data. Details of the simulation study using the nonlinear simulated data are presented in this section, while details of an analogous simulation study using linear simulated data are available in Sections 5 and 6 of the Supplementary Material (Liu and Raftery (2025)).
@@ 4.1 Data Generation
where $TN_{[0,\text{min}(X_{c,t},60)]}$ refers to the truncated normal distribution with support $[0,\text{min}(X_{c,t},60)]$. $\mathbf{X}$ and $\mathbf{Y}$ were constructed to have a generally monotonically increasing relationship similar to the relationship observed for the enrollment data. We assume that the bounds of $\mathbf{X}$ and $\mathbf{Y}$ are known to the imputer.
@@ 4.2 Analysis Model
This analysis model is uncongenial to the imputation model due to the presence of $\mathbf{Z}$, which is not included in any of the imputation models considered. We assume $\mathbf{X}$ is only available during the imputation task and is not possible to include in subsequent analyses. For example, this could occur if the refined measure $\mathbf{Y}$ comes from survey data and the coarse measure $\mathbf{X}$ comes from confidential administrative records data. Multiply imputed values of $\mathbf{Y}$ are desired for public release, but $\mathbf{X}$ cannot be made publicly available. We also assume $\mathbf{Z}$ is only available during the analysis task and is not possible to include in the imputation model. This could occur if $\mathbf{Z}$ is collected by the analyst after the multiple imputation task is complete or because $\mathbf{Z}$ is unavailable to the imputer for data confidentiality reasons. We note that omitted variable bias is likely to occur in this setting due to the omission of $\mathbf{Z}$ in the imputation model, so a data augmentation approach based on MINTS that incorporates $\mathbf{Z}$ into the imputation model is also considered in the validation exercises for comparison purposes.
Additional validation results for a random intercept analysis model can be found in Section 4 of the Supplementary Material (Liu and Raftery (2025)).
@@ 4.3 Analysis Model Validation Procedure
Data was simulated as missing following the three missing data mechanisms of Rubin (1976): Missing Completely At Random (MCAR), Missing At Random (MAR), and Missing Not at Random (MNAR). For a variable $\mathbf{X}$, MCAR occurs when $P(\mathbf{R}^{X}|\mathbf{X})=P(\mathbf{R}^{X})$. Data was simulated as missing under MCAR by assuming that each observation has the same probability of being missing. MAR occurs when $P(\mathbf{R}^{X}|\mathbf{X})=P(\mathbf{R}^{X}|\mathbf{X}_{obs})$. Data was simulated as missing under MAR by assuming that observations in earlier years are more likely to be missing. Finally, MNAR occurs when the probability of being missing depends on both the missing and the observed data and $P(\mathbf{R}^{X}|\mathbf{X})$ cannot be simplified further. Data was simulated as missing under MNAR by assuming that the probability of $\mathbf{X}$ being missing depends on the value of $\mathbf{X}$. Details of the MAR and MNAR implementations can be found in Section 3 of the Supplementary Material (Liu and Raftery (2025)).
For each missing data mechanism, data was simulated as missing at the 10%, 40%, and 80% rates. Each combination of missing data mechanism and rate was implemented simultaneously for $\mathbf{X}$ and $\mathbf{Y}$ and defines an experiment. For example, the MCAR 10% experiment corresponds to the setting where 10% of $\mathbf{X}$ was simulated as missing under MCAR and 10% of $\mathbf{Y}$ was simulated as missing under MCAR. We note that while 80% is a high rate of missingness, it is similar to the observed rate of 73.0% missingness for NER in the school enrollment data set.
All multiple imputation methods compared in the validation exercises assume that the missing data mechanism is MAR and that there is sufficient information from the observed data to adequately estimate the relationship between $\mathbf{X}$ and $\mathbf{Y}$. The experiments where data is simulated under MNAR act as a sensitivity analysis to evaluate each method’s performance when the MAR assumption is violated. Similarly, the experiments using the 80% rate of simulated missingness are considered to assess how well each method performs when there is a very large amount of missing data. In practice, substantive knowledge about the variables with missing data should be used to determine if multiple imputation is appropriate for use if the missing data mechanism is suspected to be MNAR or if there are large amounts of missing data.
@@
The performance of the point estimates $\bar{Q}_{M}$ was evaluated using the mean absolute error (MAE), calculated as $\frac{1}{N_{rep}}\sum_{r}|\bar{Q}_{M}-Q|$ where the sum is taken over all replications within each experiment. The performance of the 95% confidence intervals for the pooled estimates $\bar{Q}_{M}$ was evaluated using the mean coverage across all replications within each experiment, calculated as the proportion of intervals that contained the true value. We also evaluated the mean fraction of Fisher information about $Q$ that is missing due to nonresponse, which we abbreviate as FMI for Fraction of Missing Information. In each replication of each experiment, FMI is estimated as
@@
The mean FMI across all replications within each experiment enables assessment of the amount of information about $Q$ that is lost due to the presence of missing data (Schafer (1997a); Savalei and Rhemtulla (2012)). We note that with $M=40$ imputations, the estimates of FMI may be noisy, so in this case FMI should only be interpreted as an exploratory diagnostic (Bodner (2008); Enders (2010)).